Alyssa age is 6 years less than twice Kevin's age. Alyssa's age is 15 years minus Kevin's age. How old is Alyssa.
Answer by ikleyn(52790)

.
From the condition, you have these 2 equations


    A = 2K -6      (1)

    A = 15 - K     (2)


From this point, there are (at least) two ways to complete the solution.


Simplest way to solve this system of equations is to notice that equations (1) and (2) have identical left sides - 
hence, their right sides are equal:

    2K - 6 = 15 - K,

    2K + K = 15 + 6

    3K = 21    ==============>  K = 21/3 = 7.


So Kevin is 7 years old.  Then from equation (2)  A = 15 - 7 = 8.  

So Alyssa is 8 years old.   ANSWER

The solution is completed.